1) The following data were collected from a repeated-measures study investigating the effects of 4 treatment conditions on te using the repeated measures anova
Answer #1

We use the following model: Observation of jthe replicate of the ith treatment

= general effect + effect due to the ith treatment + random error component

= y_{ij} = \mu + \tau_i + e_{ij}

The null hypothesis is there is no significant difference between treatment or symbollically

H_0:\tau_1 = \tau_2 = \tau_3 = \tau_4

H_1: At \hspace{0.1cm} least \hspace{0.1cm} one \hspace{0.1cm} \tau_i \hspace{0.1cm} is \hspace{0.1cm} different

Sources of Variation df Sum of Squares (SS) Mean SS F Value
Treatments t - 1 SS of treatments MST MST / MSE
Error n - t SS of Error MSE
Total n - t Total SS -

Here t = 4, n = 16

Our given data

Treatments
Participants A B C D
1 6 3 3 0
2 4 4 2 2
3 4 2 0 2
4 6 3 3 0

The ANOVA Table is given by

SUMMARY
Groups Count Sum Average Variance
A 4 20 5 1.333333
B 4 12 3 0.666667
C 4 8 2 2
D 4 4 1 1.333333
ANOVA
Source of Variation SS df MS F P-value F crit
Between Groups 35 3 11.66667 8.75 0.00239 3.490295
Within Groups (error) 16 12 1.333333
Total 51 15

From the above ANOVA table we see FCalculated > FCrit

So we reject the Null hypothesis at 5% level of significance. In other words there are significant difference between treatments.

The Tukey Test for HSD is given by

HSD = \frac{M_i - M_j}{\sqrt{\frac{MS_w}{n}}}, \hspace{0.2cm}M_i>M_j

First we calculate {\sqrt{\frac{MS_w}{n}}} = {\sqrt{\frac{1.333}{4}}} = 0.57727 \approx 0.577

Now between Treatment A & B,

HSD_{AB} = \frac{5-3}{0.577} = 3.466

HSD_{AC} = \frac{5-2}{0.577} = 5.199

HSD_{AD} = \frac{5-1}{0.577} = 6.932

HSD_{BC} = \frac{3-2}{0.577} = 1.733

HSD_{BD} = \frac{3-1}{0.577} = 3.466

HSD_{CD} = \frac{2-1}{0.577} = 1.733

The Tukey's critical value at g= 0,05 is 4.08 at 15 df

So from the above calculation we conclude that Treatment A & C and Treatment A & D are significantly different

The variance % explained by the treatment effect is 35/51 = 68%

Summary

It has been observed that there is a significant difference between the treatments as observed from ANOVA table. The Treatment A and C and Treatment A & D are significantly different but there is no significant difference between B, C and D. So it may conclude that Treatment A is better than B, C and D